Two of the cylinders in an eight-cylinder car are defective and need to be replaced. If two cylinders are selected at random, what is the probability that
a. both defective cylinders are selected?
b. no defective cylinder is selected?
c. at least one defective cylinder is selected?

Answer by venugopalramana(3286)

Two of the cylinders in an eight-cylinder car are defective and need to be replaced. If two cylinders are selected at random, what is the probability that
a. both defective cylinders are selected?
number of selections of 2 cylinders from 8 = 8C2 = 8*7/2=28
ways of getting both defectives = 1
probability = 1/28
b. no defective cylinder is selected?
ways of getting 2 out of 6 good cylinders = 6C2=6*5/2=15
probability = 15/28
c. at least one defective cylinder is selected?
this is 1 - 15/28 = 13/28